Tonyo suffered a jaw injury from a road accident. Due to his jaw injury, he must wear a strap as shown. His strap produces a net upward force of 6.00N on his chin. The tension is the same throughout the strap. What is the necessary tension in the strap so that it provided the necessary upward force? The gravitational force is not included in the equation.



1
Expert's answer
2021-02-22T07:13:50-0500

Assume the angle between the parts of the straps is θ. Then according to equilibrium of forces we have


"F_\text{up}=T\text{cos}(\theta/2)+T\text{cos}(\theta/2)=2T\text{cos}(\theta/2)."

Then all we need to do is express tension:


"T=\frac{F_\text{up}}{2\text{cos}(\theta/2)}=\frac{3}{\text{cos}(\theta/2)}."

For θ=75°, the tension is 3.8 N.